GreenHab Report Apr 14th

Crew 245 GreenHab Report – 14APR2021

Green Hab Report

· Regolith Feasibility Study: regolith and Earth soil have been mixed to desired ratios in the Science Dome, and spread into the planting trays. The seeds will be added to the soil at approximately 8:30pm this evening.

· Hydroponic Garden: ongoing hydroponics experiment underway. Beakers have been set under the pipes to ameliorate spillage from leaky pipes. Intention to seal leaks with caulking tomorrow during dedicated botany time. Sealing the leak will reduce water consumption significantly.

o Each pipe currently filled with 2,000mL of water, with approx. 500mL added once today, with an estimate that the same amount will be added midday tomorrow.

· Heater: to be set to 65 degrees F.

Copyright © The Mars Society. All rights reserved. | Main Site